OpenStudy (anonymous):

A company makes similar cylindrical containers in two sizes, as shown in the table below. What is the volume, in cubic inches, of Type 1 containers? Containers Radius (in inches) Volume (in cubic inches) Type 1 3 ? Type 2 4 113.11 options: 63.62 37.70 47.72 84.83

OpenStudy (anonymous):

@psymon please help.

OpenStudy (psymon):

@babyblue123 So the question says that these cylinders are similar. What this means is that if you turn the radii of the two cylinders into a fraction, this should equal the fraction of their heights. So you have two radii, one is 3 and the other 4. So this means that if you put the height of each cylinder into a fraction, it should equal 3/4. What is given is the radius and the volume of cylinder 2. The formula for a cylinder is πr^2*h = volume. So basically we are missing the height of this second cylinder, which we can solve for. \[πr ^{2}h = volume\] r = 4 so: \[π(4)^{2}h = 113.11\] 4^2 = 16, so we will divide both sides by 16pi to get: \[h = \frac{ 113.11 }{ 16π } \approx 2.25\] Now recall that if we put the height of each cylinder into a fraction, it should = 3/4. So in order to find 3/4 of 2.25, we just multiply 2.25(3/4), which will give us approximately 1.69. This 1.69 is the height of cylinder 1. I won't solve the rest, but this is a major start. I want to make sure this makes sense or if I need to explain further.